# Clotho: Measuring Task-Specific Pre-Generation Test Adequacy for LLM Inputs

### v2

@@ 4.1 Research Questions
  • RQ1. How accurately does Clotho model the latent space of passing inputs as the reference set expands? If the modelling is accurate, the predicted likelihood of Clotho would strongly correlate with the actual passing rates of unseen inputs. We answer RQ1 primarily by comparing the Spearman rank correlation computed for Clotho and other baseline approaches that can estimate the likelihood of each input producing the correct output before generation. Further, we also investigate how much the input sampling method of Clotho affects the correlation.
  • RQ2. How effective is Clotho at prioritizing nontrivial test cases showing high failure rates for efficient testing? To answer RQ2, we predict pass/fail for inputs based on the predicted likelihood of Clotho and report ROC-AUC. We also report failure@$N$, i.e., the number of failing inputs out of top $n$ inputs when sorted according to model score: higher failure@$N$ would indicate the used score can effectively prioritise failure inducing inputs.
  • RQ3. How do Clotho’s adequacy scores complement existing post-generation uncertainty metrics? We answer RQ3 by comparing Spearman rank correlation of Clotho to post-generation measurements including internal state variance, semantic entropy, token probability, and token entropy.
  • RQ4. Do adequacy scores from Clotho transfer from open-weight models to closed-weight proprietary models as valid indicators of input difficulty? We address RQ4 by prioritizing inputs with Clotho trained on smaller open-weight LLMs and then evaluating that prioritisation on larger closed-weight models, including GPT-4o, Claude, and Gemini.
4.2 Target Tasks and Dataset
@@
  • Data Volume: We focus on datasets with sufficiently large number of inputs. First, modelling the input distribution itself requires sufficiently large number of samples, which prevents smaller benchmarks from being used. Second, we need additional inputs to check whether the learnt input distribution generalises to unseen inputs.
  • Clear Ground Truths: Many general Natural Language Understanding benchmarks, such as question answering, yield open-ended outputs that are hard to validate automatically. We instead focus on tasks with unambiguous ground truth labels.
  • Separation of Template and Inputs: Since Clotho targets specific tasks rather than open-ended dialogue, we use datasets with a clear separation between the task template (instructions) and the input (the variable part, per LLM request). This mirrors realistic use cases already supported by LLM development frameworks LangChain (2024); Sharma et al. (2025).
Table 1: Summary of Tasks, Sources, and Verification Methods Used in Our Experiments
@@
ODD-ADD 6,000 P.E. Guide AI (2025) Randomly Sampled Numbers To find sum of only odd numbers
GH-TYPO 10,000 Ours GitHub Typo Corpus Hagiwara and Mita (2019) To fix typos extracted from GitHub commits
JSON-FIX 6,563 Ours Synthetic JSON + Bugs de Jong (2022) To repair invalid JSON
MODEL-EX 9,810 P.E. Guide AI (2025) ML-ArXiv CShorten (2022), Synthetic Abstracts To extract ML model names from paper abstracts
POS-TAG 15,359 PromptPex Sharma et al. (2025); Schnabel and Neville (2024) UD_English-EWT Silveira et al. (2014) To detect Part-of-Speech tags
SPELL-CHK 10,000 Ours WordNet sentences Fellbaum (1998) + Misspellings Mitton et al. (1980) To fix misspelt words
SYN-BUG 20,518 BICS-Dataset Lee et al. (2024) Python Syntactic Bug Lee et al. (2024) To find code lines with syntactic bugs
TOPIC-CLS 7,600 PromptPex Sharma et al. (2025) AG News Zhang et al. (2015) To classify topics of news articles
The prompt templates for ODD-ADD, MODEL-EX are taken from Prompt Engineering Guide AI (2025), POS-TAG and TOPIC-CLS from PromptPex Sharma et al. (2025), and SYN-BUG from the BICS benchmark. The templates for GH-TYPO, JSON-FIX, and SPELL-CHK are written by authors to fit the corresponding tasks and datasets. For MODEL-EX, we use paper abstracts from ML-ArXiv dataset CShorten (2022) but also synthesise abstracts that use real ML model names using GPT-4o. All prompt templates, input data, as well as scripts used to synthesise inputs, are available from our replication package.
Table 2: Pass rates ($p=0.5$) of the studied open-weight LLMs on the eight tasks
@@
4.3 Models
To evaluate Clotho, we use three widely adopted Open-weight LLMs (OLMs): Llama3.1 8B Grattafiori et al. (2024), Gemma 2 9B Team et al. (2024), and Mistral 7B Jiang et al. (2023). For hidden state extraction from LIHS and LOHS, we select layers at roughly two-thirds of model depth, guided by our preliminary analysis in Section 2.1 and prior work Kossen et al. (2024); Azaria and Mitchell (2023). Table 2 reports the pass rates of these models on our tasks: an input counts as passing if it yields the correct output in a majority of generations, more than five out of ten runs. For RQ4, we additionally include three representative proprietary LLMs: GPT-4o mini [^3] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Claude Haiku [^4] https://www.anthropic.com/claude/haiku, and Gemini 2.5 Flash Lite [^5] https://deepmind.google/models/gemini/flash/. These serve as targets for testing whether adequacy scores learnt from OLMs transfer effectively to closed-weight ones.
4.4 Baselines
@@
4.4.1 Pre-generation Adequacy Metrics
@@
We first consider baselines that, like Clotho, estimate test adequacy before any output is generated: Sequence Log Likelihood (SLL), Mahalanobis Distance-based Surprise Adequacy (MDSA) Kim et al. (2020), and Base Gaussian Mixture Model (GMMb). SLL is the average log-likelihood of input tokens computed using the LLM under test. It requires no labelled data and serves as a simple, label-free adequacy baseline. MDSA computes the Mahalanobis Distance De Maesschalck et al. (2000) between a new input and the distribution of inputs in the randomly sampled reference set. By assuming a unimodal distribution, it cannot capture the more complex structures observed in input space (see Section 2). Finally, GMMb is an ablated version of Clotho that omits both the balanced sampling strategy (Section 3.1) and adaptive modelling (Section 3.2), instead fixing the number of components at five and the latent dimensionality at 10, which is the initial parameters of Clotho.
4.4.2 Post-generation Adequacy Metrics
@@
Since Clotho essentially predicts the correctness of the output induced by the given input, we can compare Clotho to post-generation uncertainty metrics: Token Probability (TOKprob) & Token-Entropy (TOKentManakul et al. (2023), Semantic Entropy (SEMentFarquhar et al. (2024), and Last Output-token Hidden State Variance (LOHSvar). TOKprob and TOKent are widely used confidence measures computed from a generated sequence: they can be computed by taking the token level likelihood and entropy at each decoding step of a single output. Following prior work, we derive sentence-level scores by averaging token log-probabilities and entropies over the sequence. SEMent requires multiple outputs from the same prompt, and computes the entropy across semantic clusters formed by the embeddings of outputs. We cluster semantically equivalent outputs among 10 generations using a DeBERTa-based NLI model He et al. (2020), before computing entropy over the resulting clusters. Finally, LOHSvar captures the variance among Last Output-token Hidden States by taking the trace of the covariance matrix of the collected LOHS vectors. This score reflects how much the model’s internal state diverge when producing different responses to the same prompt, thereby serving as a internal measure of generation variance.
4.5 Evaluation Protocols and Metrics
@@
Here we describe the process of experimentation as well as the evaluation metrics used.
4.5.1 Reference Sets and Implementation of Clotho
@@
Clotho requires a small initial reference set as a starting point. For each task, we provide 10 input-output pairs that are correct as the initial reference set: these are manually crafted and not from the datasets listed in Table 1. This is intended to simulate a realistic use case, in which a developer has a small set of working inputs for a prompt, before starting a more rigorous testing process. One exception is SYN-BUG, for which we find it difficult to craft reliably passing inputs (i.e., inputs with passing rate greater than 0.5), however simple the input is. Consequently, the initial reference set for SYN-BUG contains fewer than ten reliably passing inputs. The difficulty of this task is also noted in the original paper Lee et al. (2024). Note that the same initial reference set is used by Clotho, MDSA, and GMMb for RQ1.
@@
Clotho is implemented using GaussianMixture from sklearn.mixture with full covariance matrices Pedregosa et al. (2011). Training terminates when the log-likelihood improvement falls below $10^{-3}$ or after 500 iterations, whichever comes first. Each iteration uses a batch size of 10, and the maximum target reference set size is 500. To account for randomness in GMM training, the entire process is repeated with ten different random seeds.
4.5.2 Evaluation Metrics
@@ 5 Results
5.1 Failure Prediction Performance (RQ1)
@@
5.1.1 RQ1-1: Score Correlation with Actual Pass Rate Observed
Table 3: Comparison of pre-generation metrics across reference set sizes (Spearman rank correlation between predicted likelihoods and observed pass rates, higher values indicates better prediction performance).
Gemma Llama Mistral
@@ 5.1.1 RQ1-1: Score Correlation with Actual Pass Rate Observed
Table 3 compares the correlation between actual pass rates from ten inferences and the predicted likelihood of passing obtained from Clotho to baselines. Scores derived from distributional deviation - Clotho, MDSA, and GMMb- show weak to moderate correlations with observed pass rates, while SLL does not exhibit any meaningful correlation. Among the three distribution-based methods, Clotho achieves the highest correlation in 14 of the 24 configurations (8 tasks $\times$ 3 LLMs), while MDSA and GMMb perform best in only 3 and 7 configurations, respectively.
Figure 4: Projection of prediction scores from Clotho, MDSA, and GT pass rates (LLaMA, NN=500).
@@
Similarly, GMMb outperforms Clotho in three tasks: SPELL-CHK, GH-TYPO, and TOPIC-CLS. Note that GMMb has two major differences from Clotho: it samples inputs into the reference set randomly, and its GMM does not update component number and representation dimensionality adaptively. We note that GH-TYPO exhibits a unimodal distribution of inputs similar to that of SPELL-CHK. The structure of input distribution for TOPIC-CLS is more complex, but it is more contiguous when compared to tasks like JSON-FIX. Our conjecture is that the smaller number of mixture components used by GMMb (which is the initial value, five), as well as random sampling of inputs into the reference set, may perform reasonably well for tasks like these. Overall, the observed differences support the use of balanced sampling and adaptive modelling used by Clotho.
Figure 5: Spearman Rank Correlations from Different Sampling Strategies (NN=500)
5.1.2 RQ1-2: Comparison of Different Sampling Strategies
@@
Figure 6: Distribution of selected reference points with each sampling strategy (Llama, NN=100)
A closer examination of Figure 5 also reveals that Mistral shows a different pattern from those of Llama and Gemma: exploitation-based sampling performs poorly in general, whereas random sampling performs better than for Llama and Gemma. We conjecture that this is because Mistral exhibits substantially higher failure rates across all tasks. For example, in JSON-FIX task, only 31% of inputs achieve pass rates above 0.5, compared to 67% for Llama and 80% for Gemma. Such lower pass rate results in a larger number of smaller, more fragmented clusters of passing inputs. Such fragmented distributions are more difficult for Clotho to model, resulting in lower performance for exploitation-based sampling strategy.
Table 4: ROC-AUC scores and $p$-value from Mann-Whitney U Test pass/fail prediction ($N=500$).
@@
Gemma Llama Mistral Gemma Llama Mistral
@@
5.2 Test Prioritisation Effectiveness (RQ2)
RQ2 concerns how well Clotho can prioritise inputs that are likely to fail for more efficient testing. Table 4 shows the ROC-AUC of adequacy scores from Clotho and MDSA for the prediction of failure, as well as the $p$-values from Mann-Whitney U test of the adequacy scores between passing and failing inputs. Clotho achieves ROC-AUC score of 0.716 for unseen inputs across all three models (i.e., those not included in the reference set), when using 500 labelled references (which are, on average, 5.4% of inputs for tasks): in contrast, MDSA achieves ROC-AUC of 0.655. For all tasks and models, both Clotho and MDSA produce adequacy scores that can distinguish passing and failing inputs with statistical significance.
Table 5: Comparison of failure@$N$ for all tasks and models ($\mu$: mean, $\sigma$: standard deviation).
Gemma Llama Mistral
@@ 5.2 Test Prioritisation Effectiveness (RQ2)
Table 5 shows failure@$N$, i.e., the number of failure-inducing inputs (i.e., those whose pass rates are 0.5 or below) out of $N$ inputs selected in the descending order of adequacy metrics ($N=\{100,300,500\}$): for each pair of metric and $N$, we report average and standard deviation of failure@$N$ from 10 different orderings, each obtained from repeated runs. Note that the column Random represents random ordering and, therefore, the corresponding failure@$N$ values reflect the average ratio of failure-inducing inputs in the entire dataset for each task. The best failure@$N$ for each configuration is typeset in bold. Clotho shows strong capability in prioritizing failure-inducing inputs, with 80.4% of the top-100 ranked inputs and 75.2% of top-300 ranked inputs exposing actual failures, compared to 73.1% and 68.2% for MDSA, respectively, when aggregated over all tasks. Notably, Clotho can significantly outperform MDSA on tasks where MDSA struggles (e.g., ODD-ADD or MODEL-EX). When MDSA outperforms Clotho, the margins tend to be smaller. These results highlight the practical value of Clotho: by ranking inputs, it helps testers focus execution and labelling effort on the most challenging cases.
@@
Table 6: Comparison between Clotho and post-generation uncertainty metrics across all models and tasks.
Gemma Llama Mistral

### v3

@@ 4.1. Research Questions
  • RQ1. How accurately does Clotho model the latent space of passing inputs as the reference set expands? If the modelling is accurate, the predicted likelihood of Clotho would strongly correlate with the actual passing rates of unseen inputs. We answer RQ1 primarily by comparing the Spearman rank correlation computed for Clotho and other pre-generation baseline approaches that can estimate the likelihood of each input producing the correct output. Further, we also investigate how much the input sampling method of Clotho affects the correlation.
  • RQ2. How effective is Clotho at prioritising nontrivial test cases showing high failure rates for efficient testing? To answer RQ2, we predict pass/fail for inputs based on the predicted likelihood of Clotho and report ROC-AUC. We also report failure@$N$, i.e., the number of failing inputs out of top $n$ inputs when sorted according to model score: higher failure@$N$ would indicate the used score can effectively prioritise failure inducing inputs.
  • RQ3. How do Clotho’s adequacy scores complement existing post-generation uncertainty and confidence metrics? We answer RQ3 by comparing Spearman rank correlation of Clotho to internal state variance, semantic entropy, token probability, and token entropy.
  • RQ4. Do adequacy scores from Clotho transfer from open-weight models to closed-weight proprietary models as valid indicators of input difficulty? We address RQ4 by prioritising inputs with Clotho trained on smaller open-weight LLMs and then evaluating that prioritisation on larger closed-weight models, including GPT, Claude, and Gemini.
4.2. Target Tasks and Dataset
@@
  • Data Volume: We focus on datasets with sufficiently large number of inputs, as validating the generalisability of the input distribution modelled by Clotho to unseen test inputs requires ample data. This prevents the use of smaller benchmarks consisting of only hundreds of samples, which are generally designed for broad LLM evaluation rather than for constructing task-specific test suites.
  • Clear Ground Truths: Many general Natural Language Understanding (NLU) benchmarks, such as question answering and summarisation, yield open-ended outputs that are hard to validate both automatically and objectively. We instead focus on tasks with unambiguous ground truth labels.
  • Separation of Template and Inputs: Since Clotho targets specific tasks rather than open-ended dialogue, we use datasets with a clear separation between the task template (instructions) and the input (the variable part, per LLM request). This mirrors realistic use cases already supported by popular development frameworks for LLM applications (LangChain, 2024; Sharma et al., 2025).
Table 1. Summary of Tasks, Sources, and Verification Methods Used in Our Experiments
@@
ODD-ADD 6,000 P.E. Guide (AI, 2025) Randomly Sampled Numbers To find sum of only odd numbers
GH-TYPO 10,000 Ours GitHub Typo Corpus (Hagiwara and Mita, 2019) To fix typos extracted from GitHub commits
JSON-FIX 6,563 Ours Synthetic JSON + Bugs (de Jong, 2022) To repair invalid JSON
MODEL-EX 9,810 P.E. Guide (AI, 2025) ML-ArXiv (CShorten, 2022), Synthetic Abstracts To extract ML model names from paper abstracts
POS-TAG 15,359 PromptPex (Sharma et al., 2025; Schnabel and Neville, 2024) UD_English-EWT (Silveira et al., 2014) To detect Part-of-Speech tags
SPELL-CHK 10,000 Ours WordNet sentences (Fellbaum, 1998) + Misspellings (Mitton et al., 1980) To fix misspelt words
SYN-BUG 20,518 BICS-Dataset (Lee et al., 2024) Python Syntactic Bug (Lee et al., 2024) To find code lines with syntactic bugs
TOPIC-CLS 7,600 PromptPex (Sharma et al., 2025) AG News (Zhang et al., 2015) To classify topics of news articles
The prompt templates for ODD-ADD, MODEL-EX are taken from Prompt Engineering Guide (AI, 2025), POS-TAG and TOPIC-CLS from PromptPex (Sharma et al., 2025), and SYN-BUG from the BICS benchmark. The templates for GH-TYPO, JSON-FIX, and SPELL-CHK are written by authors to fit the corresponding tasks and datasets. For MODEL-EX, we use paper abstracts from ML-ArXiv dataset (CShorten, 2022) but also synthesise abstracts that use real ML model names using GPT-4o. All prompt templates, input data, as well as scripts used to synthesise inputs, are available from our replication package.
Table 2. Pass rates ($p=0.5$) of the studied open-weight LLMs on the eight tasks
@@
4.3. Models
To evaluate Clotho, we use three widely adopted Open-weight LLMs (OLMs): Llama3.1 8B (AI@Meta, 2024), Gemma 2 9B (Team, 2024), and Mistral 7B (Jiang et al., 2023). For hidden state extraction from LIHS and LOHS, we select layers at roughly two-thirds of model depth, guided by our preliminary analysis in Section 2.2 and prior work (Kossen et al., 2024; Azaria and Mitchell, 2023). Table 2 reports the pass rates of these models on our tasks: an input counts as passing if it yields the correct output in a majority of generations, more than five out of ten runs. For RQ4, we additionally include three representative proprietary LLMs: GPT-4o mini [^4] https://openai.com/index/gpt-4o-mini-advancing-cost-efficient-intelligence/, Claude Haiku [^5] https://www.anthropic.com/claude/haiku, and Gemini 2.5 Flash Lite [^6] https://deepmind.google/models/gemini/flash/. These serve as targets for testing whether adequacy scores learnt from OLMs transfer effectively to closed-weight ones.
4.4. Baselines
@@
4.4.1. Pre-generation Adequacy Metrics
@@
We first consider baselines that, like Clotho, estimate test adequacy before any output is generated: Sequence Log Likelihood (SLL), Mahalanobis Distance-based Surprise Adequacy (MDSA) (Kim et al., 2020), and Base Gaussian Mixture Model (GMMb). SLL is the average log-likelihood of input tokens computed using the LLM under test. It requires no labelled data and serves as a simple, label-free adequacy baseline. MDSA computes the Mahalanobis Distance (De Maesschalck et al., 2000) between a new input and the distribution of inputs in the randomly sampled reference set. By assuming a unimodal distribution, it cannot capture the more complex structures observed in input space (see Section 2). Finally, GMMb is an ablated version of Clotho that omits both the balanced sampling strategy (Section 3.1) and adaptive modelling (Section 3.2), instead fixing the number of components at five and the latent dimensionality at 10, which is the initial parameters of Clotho.
4.4.2. Post-generation Adequacy Metrics
@@
Since Clotho essentially predicts the correctness of the output induced by the given input, we can compare Clotho to post-generation uncertainty metrics: Token Probability (TOKprob) & Token-Entropy (TOKent(Manakul et al., 2023), Semantic Entropy (SEMent(Farquhar et al., 2024), and Last Output-token Hidden State Variance (LOHSvar). TOKprob and TOKent are widely used confidence measures computed from a generated sequence: they can be computed by taking the token level likelihood and entropy at each decoding step of a single output. Following prior work, we derive sentence-level scores by averaging token log-probabilities and entropies over the sequence. SEMent requires multiple outputs from the same prompt, and computes the entropy across semantic clusters formed by the embeddings of outputs. We cluster semantically equivalent outputs among 10 generations using a DeBERTa-based NLI model (He et al., 2020), before computing entropy over the resulting clusters. Finally, LOHSvar captures the variance among Last Output-token Hidden States by taking the trace of the covariance matrix of the collected LOHS vectors. This score reflects how much the model’s internal state diverge when producing different responses to the same prompt, thereby serving as an internal measure of generation variance.
4.5. Evaluation Protocols and Metrics
@@
Here, we describe the process of experimentation as well as the evaluation metrics used.
4.5.1. Reference Sets and Implementation of Clotho
@@
Clotho requires a small initial reference set as a starting point. For each task, we provide 10 inputs that mostly yield correct outputs as the initial reference set: these are manually crafted and not from the datasets listed in Table 1. This is intended to simulate a realistic use case, in which a developer has a small set of working inputs for a prompt, before starting a more rigorous testing process. One exception is SYN-BUG, for which we find it difficult to craft inputs that reliably pass with our studied OLMs (i.e., inputs with passing rate greater than 0.5), however simple the input is. Consequently, the initial reference set for SYN-BUG contains fewer than ten reliably passing inputs. The difficulty of this task is also noted in the original paper (Lee et al., 2024). Note that the same initial reference set is used by Clotho, MDSA, and GMMb for RQ1.
@@
Clotho is implemented using GaussianMixture from sklearn.mixture with full covariance matrices (Pedregosa et al., 2011). Training terminates when the log-likelihood improvement falls below $10^{-3}$ or after 500 iterations, whichever comes first. Each iteration uses a batch size of 10, and the maximum target reference set size is 500. To account for randomness in GMM training, the entire process is repeated with ten different random seeds.
4.5.2. Evaluation Metrics
@@ 5. Results
5.1. Failure Prediction Performance (RQ1)
@@
5.1.1. RQ1-1: Score Correlation with Actual Pass Rate Observed
Table 3. Comparison of pre-generation metrics across reference set sizes (Spearman rank correlation between predicted likelihoods and observed pass rates, higher values indicates better prediction performance).
Gemma 2 9B Llama 3.1 8B Mistral 7B
@@ 5.1.1. RQ1-1: Score Correlation with Actual Pass Rate Observed
Table 3 compares the correlation between actual pass rates from ten inferences and the predicted likelihood of passing obtained from Clotho to baselines. Scores derived from distributional deviation—Clotho, MDSA, and GMMb—show weak to moderate correlations with observed pass rates, while SLL does not exhibit any meaningful correlation. Among the three distribution-based methods, Clotho achieves the highest correlation in 14 of the 24 configurations (8 tasks $\times$ 3 LLMs, each with a reference set size of 500), while MDSA and GMMb perform best in only 3 and 7 configurations, respectively.
@@
Similarly, GMMb outperforms Clotho in three tasks: SPELL-CHK, GH-TYPO, and TOPIC-CLS. Note that GMMb has two major differences from Clotho: it samples inputs into the reference set randomly, and its GMM does not update component number and representation dimensionality adaptively. We note that GH-TYPO exhibits a distribution close to unimodal, similar to that of SPELL-CHK. The structure of input distribution for TOPIC-CLS is more complex, but it is more contiguous when compared to tasks like JSON-FIX. Our conjecture is that the smaller number of mixture components used by GMMb (which is the initial value, five), as well as random sampling of inputs into the reference set, may perform reasonably well for tasks like these. Overall, the observed differences support the use of balanced sampling and adaptive modelling used by Clotho.
Figure 4. Projection of prediction scores from Clotho, MDSA, and GT pass rates (Llama, NN=500).
5.1.2. RQ1-2: Comparison of Different Sampling Strategies
@@
Figure 5. Spearman Rank Correlations from Different Sampling Strategies (NN=500)
Figure 6. Distribution of selected reference points with each sampling strategy (Llama, NN=100)
A closer examination of Figure 5 also reveals that Mistral shows a different pattern from those of Llama and Gemma: exploitation-based sampling performs poorly in general, whereas random sampling performs better than for Llama and Gemma. We conjecture that this is because Mistral exhibits substantially higher failure rates across all tasks. For example, in JSON-FIX task, only 31% of inputs achieve pass rates above 0.5, compared to 67% for Llama and 80% for Gemma. Such lower pass rate results in a larger number of smaller, more fragmented clusters of passing inputs. Under limited initial data, the GMM may struggle to capture the global structure, and its entropy-based signal can become less reliable, leading to sampling that is biased toward local regions.
5.2. Test Prioritisation Effectiveness (RQ2)
RQ2 concerns how well Clotho can prioritise inputs that are likely to fail for more efficient testing. Table 4 shows the ROC-AUC of adequacy scores from Clotho and MDSA for the prediction of failure, as well as the $p$-values from Mann-Whitney U test of the adequacy scores between passing and failing inputs. Clotho achieves a ROC-AUC of 0.716 for unseen inputs across all three models (i.e., those not included in the reference set), when using 500 labelled references (which are, on average, 5.4% of inputs for tasks). In contrast, MDSA achieves a ROC-AUC of 0.655. For all tasks and models, both Clotho and MDSA produce adequacy scores that can distinguish passing and failing inputs with statistical significance.
Table 4. ROC-AUC scores and $p$-value from Mann-Whitney U Test pass/fail prediction ($N=500$).
@@ 5.2. Test Prioritisation Effectiveness (RQ2)
Gemma 2 9B Llama 3.1 8B Mistral 7B Gemma 2 9B Llama 3.1 8B Mistral 7B
@@
Table 5. Comparison of failure@$N$ for all tasks and models ($\mu$: mean, $\sigma$: standard deviation).
Gemma 2 9B Llama 3.1 8B Mistral 7B
@@
Table 5 shows failure@$N$, i.e., the number of failure-inducing inputs (i.e., those whose pass rates are 0.5 or below) out of $N$ inputs selected in the descending order of adequacy metrics ($N=\{100,300,500\}$): for each pair of metric and $N$, we report average and standard deviation of failure@$N$ from 10 different orderings, each obtained from repeated runs. Note that the column Random represents random ordering and, therefore, the corresponding failure@$N$ values reflect the average ratio of failure-inducing inputs in the entire dataset for each task. The best failure@$N$ for each configuration is typeset in bold. Clotho shows strong capability in prioritising failure-inducing inputs, with 80.4% of the top-100 ranked inputs and 75.2% of top-300 ranked inputs exposing actual failures, compared to 73.1% and 68.2% for MDSA, respectively, when aggregated over all tasks. Notably, Clotho can significantly outperform MDSA on tasks where MDSA struggles (e.g., ODD-ADD or MODEL-EX). When MDSA outperforms Clotho, the margins tend to be smaller. These results highlight the practical value of Clotho: by ranking inputs, it helps testers focus execution and labelling effort on the most challenging cases.
@@
5.3. RQ3: Complementarity Analysis
Table 6. Comparison between Clotho and post-generation uncertainty metrics across all models and tasks.
Gemma 2 9B Llama 3.1 8B Mistral 7B